# The dual complex of a $G$-variety

### v1

@@ Abstract.
equivariant birational geometry, hypersurfaces
@@ 1. Introduction
Let $G$ be a finite group which acts on a smooth projective variety $X$. A major problem in algebraic geometry is the classification of such $G$-varieties up to $G$-birational equivalence, where $X$ and $Y$ are $G$-birationally equivalent if there are invariant open sets $U\subset X$, $V\subset Y$ and an equivariant isomorphism $\varphi:U\xrightarrow{\cong}V$.
The geometry of $G$-varieties may be packaged into invariants that distinguish $G$-birational equivalence classes. This circle of ideas begins with the observation that the presence of an $H$-fixed point is a $G$-birational invariant of $X$ for $H\subset G$ abelian, over an algebraically closed field of characteristic zero [15, App. A]. The recent introduction of equivariant Burnside groups [11] has led to many new advances in this area. The $G$-birational invariants which take values in these groups incorporate the birational types of fixed loci of all abelian subgroups $H\subset G$, actions of $H$ on the normal spaces to these fixed loci, and residual actions on their orbits. They have found numerous applications in classifying $G$-varieties of low dimension, such as projective representations of finite groups [12, 18], algebraic tori [13], and actions on other rationally connected threefolds [5].
@@
Existing $G$-birational invariants do not directly measure how components of the non-free locus of a $G$-variety intersect. Therefore, in this paper, we introduce a notion of dual complexes for $G$-varieties, which record intersection data for the part of the non-free locus that “looks most like the boundary of a toric variety”. We show that the top homology group of this complex is a $G$-birational invariant (Theorem 2.4) and use this result to find new examples of non-birational actions. In particular, we prove that actions of rank $n$ abelian subgroups $G\subset\mathrm{Aut}(\mathbb{P}^{n+1})$ on invariant smooth hypersurfaces $X\subset\mathbb{P}^{n+1}$ of degree $d\geq 3$ are never $G$-birational to a toric action on a smooth toric variety, except possibly in the case of cubic surfaces (Theorem 3.2).
@@ 2. Dual Complexes
Suppose $X$ is a smooth projective $G$-variety with abelian stabilizers. By Luna’s slice theorem [14, Lemme III.1], for any closed point $p\in X$, there is an invariant affine open neighborhood $U$ of $p$ and a $\mathrm{stab}_{G}(p)$-equivariant morphism $\varphi:U\rightarrow T_{p}X$ such that $\varphi$ is étale at $p$ and $\varphi(p)=0$. In other words, the action is “étale-locally linear”. Therefore, for any subgroup $H\subset\mathrm{stab}_{G}(p)$, $T_{p}X^{H}\cong(T_{p}X)^{H}$ and $X^{H}$ is smooth at $p$ (though globally $X^{H}$ may be a union of connected components of different dimensions). It follows that the representation of $\mathrm{stab}_{G}(p)$ on $T_{p}X$ is faithful. Thus, for any abelian group $H\subset G$, any connected component of $X^{H}$ has codimension at least $\mathrm{rank}(H)$ in $X$.
@@
is a nontrivial cyclic group for each $i$. The subgroup $H_{i}$ is the stabilizer of a hyperplane in $T_{p}X$, so this hyperplane is the tangent space to a divisor $D_{i}$ in $X^{H_{i}}$ passing through $p$. The divisors $D_{1},\ldots,D_{k}$ have simple normal crossings at $p$ since all of their intersections are smooth of the expected dimension near $p$ [17, Lemma 0BIA]. ∎
@@
By the $G$-equivariant weak factorization theorem [1], any birational map of smooth projective $G$-varieties may be decomposed into a sequence of blowups and blowdowns in smooth centers. If the varieties both have abelian stabilizers, we may guarantee that the intermediate varieties have abelian stabilizers as well (cf. [11, Proposition 3.6]). {commentA} Indeed, that proposition shows that any two varieties $X$ and $X^{\prime}$ satisfying Assumption 2 of [11] are connected by a series of blowups and blowdowns in smooth centers, where each intermediate variety satisfies Assumption 2 as well. One of the conditions of Assumption 2 is having abelian stabilizers. Though this isn’t sufficient to satisfy Assumption 2, there are equivariant blowups of $X$ and $X^{\prime}$ satisfying Assumption 2 [11, Proposition 3.3], and these can be connected by intermediate varieties satisfying the assumption. Since a blowup of a variety with abelian stabilizers has abelian stabilizers, this completes the argument. Therefore, it suffices to show that the homology group $H_{n-1}$ of the dual complex is unchanged by a single blowup $X^{\prime}\coloneqq\mathrm{Bl}_{W}(X)\rightarrow X$ of a variety $X$ of dimension $n$ with abelian stabilizers in a $G$-stable center $W$, which is a disjoint union of smooth irreducible varieties. We may assume that $G$ acts transitively on the connected components of $W$.
@@
To see that the effect of the blowup is a stellar subdivision, let $Z$ be any connected component of $W$, and $D_{1},\ldots,D_{k}$ the divisors from Lemma 2.1 containing $Z$. Each $D_{i}$ is fixed by a nontrivial cyclic subgroup $H_{i}\subset G$ and the sum of these subgroups has rank $k$. Denote by $E_{Z}$ the exceptional divisor over $Z$. Any maximal rank stratum contained in $Z$ is a component of an intersection of the form $D_{1}\cap\cdots\cap D_{k}\cap D_{1}^{\prime}\cap\cdots\cap D_{\ell}^{\prime}$ for some additional divisors $D_{1}^{\prime},\ldots,D_{\ell}^{\prime}$, where the stabilizer of a general point on each is a nontrivial cyclic group. Furthermore, these divisors have simple normal crossings in a neighborhood of the intersection. It’s well known that the dual complex of the blow up of a stratum in an snc intersection is the stellar subdivision of the original complex (see, e.g., [2, Section 7]). It remains to see that all the strata of the blowup “count” for our notion of dual complex, namely that for each proper subset $I\subset\{1,\ldots,k\}$, $E_{Z}\cap D_{I}\cap D_{1}^{\prime}\cap\cdots\cap D_{\ell}^{\prime}$ is a maximal rank stratum in $X^{\prime}$, where $D_{I}=\bigcap_{i\in I}D_{i}$.
@@
Let $H={\mathbb{Z}}/d_{1}\oplus\cdots\oplus{\mathbb{Z}}/d_{n}$ be a rank $n$ abelian group acting diagonally on $\mathbb{A}^{n}$. Then any $H$-stable smooth subvariety $Y\subset\mathbb{A}^{n}$ through $0\in\mathbb{A}^{n}$ is a coordinate plane.
@@
Since $Y$ is smooth and passes through 00, its tangent cone is some linear subspace of $\mathbb{A}^{n}$. It is also $H$-stable, so we may assume the initial ideal of $I$ (which cuts out the tangent cone) is $\mathrm{in}(I)=(x_{1},\ldots,x_{r})$, where $\dim(Y)=n-r$. Among the linear forms in $\mathrm{in}(I)$, only $x_{1},\ldots,x_{r}$ are preserved up to scaling by $H$, so the generating set contains an element of the form $x_{1}+h_{1}$, where $h_{1}$ only has terms of degree at least $2$. Since $x_{1}$ and $h_{1}$ must have the same $G$-degree, it follows that $x_{1}$ divides $h_{1}$. Since $I$ is prime, $x_{1}\in I$. Similarly, $x_{2},\ldots,x_{r}\in I$. For dimension reasons, $I=(x_{1},\ldots,x_{r})$, so $Y$ is a coordinate plane. ∎
Returning to the proof of Theorem 2.4, we claim that the assumption on the stabilizers of points of $W$ means that $W$ does not contain a maximal rank stratum of dimension 00. Suppose by way of contradiction that $W$ does contain such a point $p$ with stabilizer $H$, with $\mathrm{rank}(H)=n$. There is a neighborhood $U$ of $p$ and an $H$-equivariant morphism $\varphi:U\rightarrow T_{p}X$ which is étale at $p$ and sends $p$ to 00, by Luna’s slice theorem. Restricting to a smaller open $V\subset U$ containing $p$ where $\varphi$ is étale, we have that $\phi(V\cap W)$ is an $H$-stable closed subvariety of an open neighborhood of 00 in $T_{p}X\cong\mathbb{A}^{n}$ which passes through 00 and is smooth there. The group $H$ is rank $n$ and acts linearly on $T_{p}X$. After diagonalizing this action, we may apply Lemma 2.5 to conclude that $\varphi(V\cap W)$ coincides with a coordinate plane near 00 in $T_{p}X$. Every coordinate plane is the image of a maximal rank stratum, so $W$ and some stratum $Z$ coincide étale-locally near $x$. Hence $W$ is a union of maximal rank strata, a contradiction.
@@ 3. Applications
Let $X$ be a smooth projective toric variety of dimension $n$ and $G\subset\mathbb{G}_{\operatorname{m}}^{n}$ a full rank finite subgroup of the torus. The non-free locus for the $G$-action is precisely the toric boundary, and each stratum of the boundary has maximal rank by the assumption that $G$ has rank $n$. {commentA} For instance, consider any toric divisor of $X$, corresponding to a ray with primitive point $r=(c_{1},\ldots,c_{n})$ in the lattice $N$ of one-parameter subgroups of $\mathbb{G}_{\operatorname{m}}^{n}$. Our claim amounts to the fact that any such one-parameter subgroup intersects $G$ nontrivially. The group fixing $D_{r}$ for $r$ as above is $\{(t^{c_{1}},\ldots,t^{c_{n}}):t\in{\mathbb{C}}^{*}\}$. Suppose $G$ has elementary abelian subgroup $({\mathbb{Z}}/p)^{\oplus n}$. Then plugging in $t=\zeta_{p}$, we find a nontrivial element $(\zeta_{p}^{c_{1}},\ldots,\zeta_{p}^{c_{n}})$ in the intersection of $G$ with the one-parameter subgroup (it is nontrivial because not all $c_{i}$ can be divisible by $p$). The rays spanning any cone in the fan of $X$ form part of a ${\mathbb{Z}}$-basis for $N$, so we can use a similar argument to find a subgroup of the appropriate rank fixing the corresponding troric stratum. Since $X$ is smooth and projective, its fan is the span of the faces of some simplicial polytope $P$ in the lattice $N$ of one-parameter subgroups [9, Theorem VII.3.11]. Therefore, the dual complex $\mathcal{D}_{G}(X)$ is simply $P$. In particular, $\mathcal{D}_{G}(X)$ is homeomorphic to the sphere $S^{n-1}$, so the top homology of $\mathcal{D}_{G}(X)$ is $H_{n-1}(S^{n-1})\cong{\mathbb{Z}}$.
@@
Let $n$ and $d$ be positive integers such that $n\geq 1$, $d\geq 3$, and $(n,d)\neq(2,3)$. Then for any rank $n$ abelian subgroup $G\subset\mathrm{Aut}(\mathbb{P}^{n+1})$ and any $G$-invariant smooth hypersurface $X\subset\mathbb{P}^{n+1}$ of degree $d$, $X$ is not $G$-birational to an action by a subgroup of the torus on a smooth projective toric variety. In particular, the action is not linearizable.
@@
We claim that two of $\chi_{\mathrm{triv}},\chi_{1},\ldots,\chi_{n+1}$ must be the same, as long as $n\geq 3$. (The theorem automatically holds for curves of degree $d\geq 3$ and surfaces of degree $d\geq 4$ because these are all irrational, so we ignore these cases.) Indeed, if the characters are all distinct, the only $G$-fixed points on $\mathbb{P}^{n+1}$ are the coordinate points. We may assume $X$ contains the coordinate point $p$ of $x_{0}$ (which $G$ acts on by the trivial character). Then $T_{p}X\subset T_{p}\mathbb{P}^{n+1}$ is a sub-$G$-representation. Since $T_{p}\mathbb{P}^{n+1}$ is a sum of distinct characters, $T_{p}X$ coincides with the tangent space to a coordinate hyperplane, which we can assume is $\{x_{1}=0\}$ after relabeling. It also follows that $T_{p}X\cong\chi_{2}\oplus\cdots\oplus\chi_{n+1}$ is faithful, so $\chi_{2},\ldots,\chi_{n+1}$ form a basis for $G^{*}\cong\mathbb{F}_{p}^{n}$. By Lemma 2.1, $p$ must be the intersection of $n$ divisors on $X$ which are each fixed by a nontrivial subgroup of $G$. These divisors are precisely $\{x_{2}=0\}\cap X,\ldots,\{x_{n+1}=0\}\cap X$. Indeed, only coordinate strata in $\mathbb{P}^{n+1}$ are fixed by a nontrivial subgroup, and $X$ cannot contain any codimension $2$ coordinate plane of $\mathbb{P}^{n+1}$ when it is smooth of degree $d>1$ and $n\geq 3$ [8, Corollary 6.26].
@@
We may therefore suppose that the $G$-action on $V\cong k^{n+2}$ has the following form: each summand of $({\mathbb{Z}}/p)^{\oplus n}$ acts by multiplication by $p$th roots of unity on a distinct coordinate $x_{2},\ldots,x_{n+1}$, and the action on the first two coordinates is trivial. Since $X$ cannot contain a codimension $2$ coordinate plane of $\mathbb{P}^{n+1}$, the only divisors fixed by nontrivial subgroups in $X$ are the intersections $D_{i}\coloneqq X\cap\{x_{i}=0\}$, $i=2,\ldots,{n+1}$, where $D_{i}$ is fixed by the $i$th summand ${\mathbb{Z}}/p$ of $G$. Every intersection of $k<n$ of these divisors has a rank $k$ stabilizer, so it is smooth and irreducible of dimension $n-k\geq 1$. Thus, the $(n-2)$-skeleton of $\mathcal{D}_{G}(X)$ is the same as the $(n-2)$-skeleton of the standard $(n-1)$-simplex. However, $D_{2}\cap\cdots\cap D_{n+1}$ is a union of $d$ distinct points (or else the divisors do not intersect transversely), so $\mathcal{D}_{G}(X)$ has $d$ top-dimensional simplices with identical boundary. Therefore, $H_{n-1}(\mathcal{D}_{G}(X))\cong{\mathbb{Z}}^{d-1}$. By Theorem 2.4 and Example 3.1, $X$ is not $G$-birational to a smooth toric variety with action by a full rank finite subgroup of the torus whenever $d\geq 3$. This implies in particular that the $G$-action on $X$ is not linearizable. ∎
@@
Theorem 3.2 does not hold for the case of quadrics, where $d=2$, and for cubic surfaces, where $(n,d)=(2,3)$. For quadrics, any $G$-action with a $G$-fixed point $p$ is linearizable, becuase the projection from $p$ gives a $G$-equivariant birational map to projective space. Of course, an action by an abelian $G$ without a fixed point is not linearizable, since linearizable actions have fixed points.
There is also a counterexample for smooth cubic surfaces, which are always rational. Let $X=\{x_{0}^{3}+x_{1}^{3}+x_{2}^{3}+x_{3}^{3}=0\}\subset\mathbb{P}^{3}$ be the Fermat cubic surface and $G\subset\mathrm{Aut}(X)$ the normal Klein four-subgroup of the $S_{4}$ permuting the four coordinates. Each involution in $G$ fixes a line and three isolated points in $X$ (these involutions are of type 2B in the classification of Dolgachev-Duncan [7]). No nontrivial element of $G$ fixes a positive genus curve, so $G$ is conjugate to a subgroup of $\mathrm{Aut}(\mathbb{P}^{1}\times\mathbb{P}^{1})$ or $\mathrm{Aut}(\mathbb{P}^{2})$ in the plane Cremona group [3, Theorem 5]. Since $G$ also fixes a point, the subgroup must be linearizable.
@@
There is an alternate approach to the second part of the proof of Theorem 3.2 suggested by Yuri Tschinkel, using the formalism of incompressible symbols in equivariant Burnside groups. We sketch the argument below in the case of $G\coloneqq({\mathbb{Z}}/d)^{\oplus n}$ acting on the first $n$ coordinates of a hypersurface $X$ of dimension $n$ and degree $d$.

### v2

@@ Abstract.
equivariant birational geometry, hypersurfaces, linearizable actions
@@ 1. Introduction
Let $G$ be a finite group which acts on a smooth projective variety $X$. A major problem in algebraic geometry is the classification of such $G$-varieties up to $G$-birational equivalence, where $X$ and $Y$ are $G$-birationally equivalent if there are invariant open sets $U\subset X$, $V\subset Y$ and a $G$-equivariant isomorphism $\varphi:U\xrightarrow{\cong}V$.
The geometry of $G$-varieties may be packaged into invariants that distinguish $G$-birational equivalence classes. This circle of ideas begins with the observation that the presence of an $H$-fixed point is a $G$-birational invariant of $X$ for $H\subset G$ abelian, over an algebraically closed field of characteristic zero [18, App. A]. The recent introduction of equivariant Burnside groups [14] has led to many new advances in this area. The $G$-birational invariants which take values in these groups incorporate the birational types of fixed loci of all abelian subgroups $H\subset G$, actions of $H$ on the normal spaces to these fixed loci, and residual actions on their orbits. They have found numerous applications in classifying $G$-varieties of low dimension, such as projective representations of finite groups [15, 21], algebraic tori [16], and actions on other rationally connected threefolds [6].
@@
Existing $G$-birational invariants do not directly measure how components of the non-free locus of a $G$-variety intersect. Therefore, in this paper, we introduce a notion of dual complexes for $G$-varieties, which roughly records intersection data for the part of the non-free locus that is an snc divisor. We show that the top homology group of this complex is a $G$-birational invariant (Theorem 2.4) and use this result to find new examples of non-birational actions. In particular, we prove that actions of rank $n$ abelian subgroups $G\subset\mathrm{Aut}(\mathbb{P}^{n+1})$ on invariant smooth hypersurfaces $X\subset\mathbb{P}^{n+1}$ of degree $d\geq 3$ are never $G$-birational to a toric action on a smooth toric variety, apart from some exceptional actions on cubic surfaces (Theorem 3.2).
@@ 2. Dual Complexes
Suppose $X$ is a smooth projective $G$-variety with abelian stabilizers. By Luna’s slice theorem [17, Lemme III.1], for any closed point $p\in X$, there is an invariant affine open neighborhood $U$ of $p$ and a $\mathrm{stab}_{G}(p)$-equivariant morphism $\varphi:U\rightarrow T_{p}X$ such that $\varphi$ is étale at $p$ and $\varphi(p)=0$. In other words, the action is “étale-locally linear”. Therefore, for any subgroup $H\subset\mathrm{stab}_{G}(p)$, $T_{p}X^{H}\cong(T_{p}X)^{H}$ and $X^{H}$ is smooth at $p$ (though globally $X^{H}$ may be a union of connected components of different dimensions). It follows that the representation of $\mathrm{stab}_{G}(p)$ on $T_{p}X$ is faithful. Thus, for any abelian group $H\subset G$, any connected component of $X^{H}$ has codimension at least $\mathrm{rank}(H)$ in $X$.
@@
is a nontrivial cyclic group for each $i$. The subgroup $H_{i}$ is the stabilizer of a hyperplane in $T_{p}X$, so this hyperplane is the tangent space to a divisor $D_{i}$ in $X^{H_{i}}$ passing through $p$. The divisors $D_{1},\ldots,D_{k}$ have simple normal crossings at $p$ since all of their intersections are smooth of the expected dimension near $p$ [20, Lemma 0BIA]. ∎
@@
By the $G$-equivariant weak factorization theorem [1], any birational map of smooth projective $G$-varieties may be decomposed into a sequence of blowups and blowdowns in smooth centers. If the varieties both have abelian stabilizers, we may guarantee that the intermediate varieties have abelian stabilizers as well (cf. [14, Proposition 3.6]). {commentA} Indeed, that proposition shows that any two varieties $X$ and $X^{\prime}$ satisfying Assumption 2 of [14] are connected by a series of blowups and blowdowns in smooth centers, where each intermediate variety satisfies Assumption 2 as well. One of the conditions of Assumption 2 is having abelian stabilizers. Though this isn’t sufficient to satisfy Assumption 2, there are equivariant blowups of $X$ and $X^{\prime}$ satisfying Assumption 2 [14, Proposition 3.3], and these can be connected by intermediate varieties satisfying the assumption. Since a blowup of a variety with abelian stabilizers has abelian stabilizers, this completes the argument. Therefore, it suffices to show that the homology group $H_{n-1}$ of the dual complex is unchanged by a single blowup $X^{\prime}\coloneqq\mathrm{Bl}_{W}(X)\rightarrow X$ of a variety $X$ of dimension $n$ with abelian stabilizers in a $G$-stable center $W$, which is a disjoint union of smooth irreducible varieties. We may assume that $G$ acts transitively on the connected components of $W$.
@@
To see that the effect of the blowup is a stellar subdivision, let $Z$ be any connected component of $W$, and $D_{1},\ldots,D_{k}$ the divisors from Lemma 2.1 containing $Z$. We continue to use the notations $H_{i}$ and $H$ from the lemma for the groups fixing $D_{i}$ and $Z$ pointwise, respectively. So each $D_{i}$ is fixed by a nontrivial cyclic subgroup $H_{i}\subset G$ and the sum $H$ of these subgroups has rank $k$. Denote by $E_{Z}$ the exceptional divisor over $Z$. Any maximal rank stratum contained in $Z$ is a component of an intersection of the form $D_{1}\cap\cdots\cap D_{k}\cap D_{1}^{\prime}\cap\cdots\cap D_{\ell}^{\prime}$ for some additional divisors $D_{1}^{\prime},\ldots,D_{\ell}^{\prime}$, where the stabilizer of a general point on each is a nontrivial cyclic group. Furthermore, these divisors have simple normal crossings in a neighborhood of the intersection. It’s well known that the dual complex of the blow up of a stratum in an snc intersection is the stellar subdivision of the original complex in the simplex corresponding to that stratum (see, e.g., [2, Section 7]). It remains to see that all the strata of the blowup “count” for our notion of the dual complex, namely that for each proper subset $I\subset\{1,\ldots,k\}$, $E_{Z}\cap D_{I}\cap D_{1}^{\prime}\cap\cdots\cap D_{\ell}^{\prime}$ is a maximal rank stratum in $X^{\prime}$, where $D_{I}=\bigcap_{i\in I}D_{i}$.
@@
Let $H={\mathbb{Z}}/d_{1}\oplus\cdots\oplus{\mathbb{Z}}/d_{n}$ be a rank $n$ abelian group acting diagonally on $\mathbb{A}^{n}$. Then any $H$-stable smooth subvariety $Y\subset\mathbb{A}^{n}$ through $0\in\mathbb{A}^{n}$ is a coordinate subspace.
@@
Since $Y$ is smooth and passes through 00, its tangent cone is some linear subspace of $\mathbb{A}^{n}$. It is also $H$-stable, so we may assume the initial ideal of $I$ (which cuts out the tangent cone) is $\mathrm{in}(I)=(x_{1},\ldots,x_{r})$, where $\dim(Y)=n-r$. Among the linear forms in $\mathrm{in}(I)$, only $x_{1},\ldots,x_{r}$ are preserved up to scaling by $H$, so the generating set contains an element of the form $x_{1}+h_{1}$, where $h_{1}$ only has terms of degree at least $2$. Since $x_{1}$ and $h_{1}$ must have the same $H$-degree, it follows that $x_{1}$ divides $h_{1}$. Since $I$ is prime, $x_{1}\in I$. Similarly, $x_{2},\ldots,x_{r}\in I$. For dimension reasons, $I=(x_{1},\ldots,x_{r})$, so $Y$ is a coordinate plane. ∎
Returning to the proof of Theorem 2.4, we claim that the assumption on the stabilizers of points of $W$ means that $W$ does not contain a maximal rank stratum of dimension 00. Suppose by way of contradiction that $W$ does contain such a point $p$ with stabilizer $H$, with $\mathrm{rank}(H)=n$. There is a neighborhood $U$ of $p$ and an $H$-equivariant morphism $\varphi:U\rightarrow T_{p}X$ which is étale at $p$ and sends $p$ to 00, by Luna’s slice theorem. Restricting to a smaller open $V\subset U$ containing $p$ where $\varphi$ is étale, we have that $\varphi(V\cap W)$ is an $H$-stable closed subvariety of an open neighborhood of 00 in $T_{p}X\cong\mathbb{A}^{n}$ which passes through 00 and is smooth there. The group $H$ is rank $n$ and acts linearly on $T_{p}X$. After diagonalizing this action, we may apply Lemma 2.5 to conclude that $\varphi(V\cap W)$ coincides with a coordinate plane near 00 in $T_{p}X$. Every coordinate plane is the image of a maximal rank stratum, so $W$ and some stratum $Z$ coincide étale-locally near $x$. Hence $W$ is a union of maximal rank strata, a contradiction.
@@ 3. Applications
Let $X$ be a smooth projective toric variety of dimension $n$ and $G\subset\mathbb{G}_{\operatorname{m}}^{n}$ a full rank finite subgroup of the torus. The non-free locus for the $G$-action is precisely the toric boundary, and each stratum of the boundary has maximal rank by the assumption that $G$ has rank $n$. {commentA} For instance, consider any toric divisor of $X$, corresponding to a ray with primitive point $r=(c_{1},\ldots,c_{n})$ in the lattice $N$ of one-parameter subgroups of $\mathbb{G}_{\operatorname{m}}^{n}$. Our claim amounts to the fact that any such one-parameter subgroup intersects $G$ nontrivially. The group fixing $D_{r}$ for $r$ as above is $\{(t^{c_{1}},\ldots,t^{c_{n}}):t\in{\mathbb{C}}^{*}\}$. Suppose $G$ has elementary abelian subgroup $({\mathbb{Z}}/p)^{\oplus n}$. Then plugging in $t=\zeta_{p}$, we find a nontrivial element $(\zeta_{p}^{c_{1}},\ldots,\zeta_{p}^{c_{n}})$ in the intersection of $G$ with the one-parameter subgroup (it is nontrivial because not all $c_{i}$ can be divisible by $p$). The rays spanning any cone in the fan of $X$ form part of a ${\mathbb{Z}}$-basis for $N$, so we can use a similar argument to find a subgroup of the appropriate rank fixing the corresponding troric stratum. Since $X$ is smooth and projective, its fan is the span of the faces of some simplicial polytope $P$ in the lattice $N$ of one-parameter subgroups [10, Theorem VII.3.11]. Therefore, the dual complex $\mathcal{D}_{G}(X)$ is simply $P$. In particular, $\mathcal{D}_{G}(X)$ is homeomorphic to the sphere $S^{n-1}$, so the top homology of $\mathcal{D}_{G}(X)$ is $H_{n-1}(S^{n-1})\cong{\mathbb{Z}}$.
@@
Let $n$ and $d$ be positive integers such that $n\geq 1$, $d\geq 3$, and $(n,d)\neq(2,3)$. Then the action of any rank $n$ abelian group $G$ on a smooth hypersurface $X\subset\mathbb{P}^{n+1}$ of degree $d$ is not $G$-birational to an action by a subgroup of the torus on a smooth projective toric variety. In particular, the action is not linearizable.
@@
Smooth plane curves of degree at least $3$ and smooth surfaces of degree at least $4$ in $\mathbb{P}^{3}$ are not rational, so we may immediately assume that $n\geq 3$. In that setting, the Grothendieck-Lefschetz theorem [11, Exposé XII, Corollaire 3.6] implies that the group $G\subset\mathrm{Aut}(X)$ lifts to a subgroup $G\subset\mathrm{Aut}(\mathbb{P}^{n+1})\cong\mathrm{PGL}_{n+2}(k)$.
@@
We claim that two of $\chi_{\mathrm{triv}},\chi_{1},\ldots,\chi_{n+1}$ must be the same. Indeed, if the characters are all distinct, the only $G$-fixed points on $\mathbb{P}^{n+1}$ are the coordinate points. We may assume $X$ contains the coordinate point $p=[1:0:\cdots:0]$ and that $G$ acts on the corresponding line by the trivial character. Then $T_{p}X\subset T_{p}\mathbb{P}^{n+1}$ is a sub-$G$-representation. Since $T_{p}\mathbb{P}^{n+1}$ is a sum of distinct characters, $T_{p}X$ coincides with the tangent space to a coordinate hyperplane, which we can assume is $\{x_{1}=0\}$ after relabeling. It also follows that $T_{p}X\cong\chi_{2}\oplus\cdots\oplus\chi_{n+1}$ is faithful, so $\chi_{2},\ldots,\chi_{n+1}$ form a basis for $G^{*}\cong\mathbb{F}_{p}^{n}$. By Lemma 2.1, $p$ must be the intersection of $n$ divisors on $X$ which are each fixed by a nontrivial subgroup of $G$. These divisors are precisely $\{x_{2}=0\}\cap X,\ldots,\{x_{n+1}=0\}\cap X$. Indeed, only coordinate strata in $\mathbb{P}^{n+1}$ are fixed by a nontrivial subgroup, and $X$ cannot contain any codimension $2$ coordinate plane of $\mathbb{P}^{n+1}$ when it is smooth of degree $d>1$ and $n\geq 3$ [9, Corollary 6.26].
@@
We may therefore assume that the $G$-action on $V\cong k^{n+2}$ has the following form: each summand of $({\mathbb{Z}}/p)^{\oplus n}$ acts by multiplication by $p$th roots of unity on a distinct coordinate $x_{2},\ldots,x_{n+1}$, and the action on the first two coordinates is trivial. Since $X$ cannot contain a codimension $2$ coordinate plane of $\mathbb{P}^{n+1}$, the only divisors fixed by nontrivial subgroups in $X$ are the intersections $D_{i}\coloneqq X\cap\{x_{i}=0\}$, $i=2,\ldots,{n+1}$, where $D_{i}$ is fixed by the $i$th summand ${\mathbb{Z}}/p$ of $G$. Every intersection of $k<n$ of these divisors has a rank $k$ stabilizer, so it is smooth and irreducible of dimension $n-k\geq 1$. Thus, the $(n-2)$-skeleton of $\mathcal{D}_{G}(X)$ is the same as the $(n-2)$-skeleton of the standard $(n-1)$-simplex. However, $D_{2}\cap\cdots\cap D_{n+1}$ is a union of $d$ distinct points (or else the divisors do not intersect transversely), so $\mathcal{D}_{G}(X)$ has $d$ top-dimensional simplices with identical boundary. Therefore, $H_{n-1}(\mathcal{D}_{G}(X))\cong{\mathbb{Z}}^{d-1}$. By Theorem 2.4 and Example 3.1, $X$ is not $G$-birational to a smooth toric variety with action by a full rank finite subgroup of the torus whenever $d\geq 3$. This implies in particular that the $G$-action on $X$ is not linearizable. ∎
{commentA}
Remark 3.3.
A related question asks whether any action of $G$ on a hypersurface $X$ can be projectively linearizable, i.e., $G$-equivariantly isomorphic to any $G$ action on $\mathbb{P}^{n}$, not necessarily linear. The techniques of Theorem 3.2 do not directly answer this question, but we note that projectively linearizable and linearizable are frequently equivalent for a fixed pair $(n,d)$. In particular, we claim that they are equivalent when $n\geq 3$, $d\geq 3$, and $\gcd(n+2,d)=1$.
Indeed, given the finite group $G\subset\mathrm{Aut}(X)$, we know $G$ lifts to $\mathrm{Aut}(\mathbb{P}^{n+1})$. But the group $G$ thus also acts on $K_{\mathbb{P}^{4}}$ and $K_{X}$. These two line bundles are $\mathcal{O}_{X}(-n-2)$ and $\mathcal{O}_{X}(d-n-2)$, respectively. If $\gcd(d,n+2)=1$, then some product of tensor powers of these is $\mathcal{O}_{X}(1)$, so we get a $G$-action on $H^{0}(X,\mathcal{O}_{X}(1))$ inducing the action on $\mathbb{P}^{n+1}$, meaning that the $G$-action is linear. This in turn means that the Amitsur group $\mathrm{Am}(X,G)$ vanishes [4, Section 6]. This group is also an equivariant birational invariant, so if the action were projectively linearizable, the corresponding action of $G$ on $\mathbb{P}^{n}$ would also have vanishing Amitsur group. Hence it would be linear. It follows that projective linearizability implies linearizability in this case.
@@
Theorem 3.2 does not hold for the case of quadrics, where $d=2$, and for cubic surfaces, where $(n,d)=(2,3)$. For quadrics, any $G$-action with a $G$-fixed point $p$ is linearizable, because the projection from $p$ gives a $G$-equivariant birational map to projective space. Of course, an action by an abelian group $G$ without a fixed point is not linearizable, since linearizable actions have fixed points.
There is also a counterexample for smooth cubic surfaces, which are always rational. Let $X=\{x_{0}^{3}+x_{1}^{3}+x_{2}^{3}+x_{3}^{3}=0\}\subset\mathbb{P}^{3}$ be the Fermat cubic surface and $G\subset\mathrm{Aut}(X)$ the normal Klein four-subgroup of the $S_{4}$ permuting the four coordinates. Each involution in $G$ fixes a line and three isolated points in $X$ (these involutions are of type 2B in the classification of Dolgachev-Duncan [8]). No nontrivial element of $G$ fixes a positive genus curve, so $G$ is conjugate to a subgroup of $\mathrm{Aut}(\mathbb{P}^{1}\times\mathbb{P}^{1})$ or $\mathrm{Aut}(\mathbb{P}^{2})$ in the plane Cremona group [3, Theorem 5]. Since $G$ also fixes the point $(1:-1:1:-1)\in X$, the subgroup must be linearizable.
Example 3.5 (Rational Hypersurfaces).
For every even $n\geq 2$, the dual complex can be used to identify nonlinearizable actions on rational cubic hypersurfaces of dimension $n$. For instance, let $X$ be the Fermat cubic hypersurface
$X=\{x_{0}^{3}+\cdots+x_{n+1}^{3}=0\}\subset\mathbb{P}^{n+1}.$
This hypersurface contains a pair of complementary linear subspaces of dimension $n/2$, so it is rational [12, Corollary 1.5.11]. However, there are many actions of the group $G\coloneqq({\mathbb{Z}}/3)^{\oplus n}$ on this hypersurface, such as the action by $3$rd roots of unity on the first $n$ coordinates. By Theorem 3.2, this action is not linearizable. It is not even linearizable in dimension $2$, since the dual complex is homeomorphic to $S^{1}\vee S^{1}$. In combination with Example 3.4, this gives an example of a single surface with two actions by rank $2$ abelian groups such that each has a fixed point, but one is linearizable and one is not.
@@
An alternate approach to part of the proof of Theorem 3.2 was later suggested by Yuri Tschinkel, using the formalism of incompressible symbols in equivariant Burnside groups. We sketch the argument below in the case of $G\coloneqq({\mathbb{Z}}/d)^{\oplus n}$ acting on the first $n$ coordinates of a hypersurface $X$ of dimension $n$ and degree $d$.